# CodeARC: Benchmarking Reasoning Capabilities of LLM Agents for Inductive Program Synthesis

### v1

@@ 5.5 Case Study
Figure 4 shows an interaction trace from our benchmark. The model starts by querying the ground-truth function with edge-case inputs, aiming to probe its behavior beyond the initial examples. It then synthesizes a candidate solution using pairwise comparisons, which passes the given examples but fails on a counterexample with unhashable elements. From the error message, the model correctly infers that the ground-truth function raises a TypeError, while its own does not. On the second attempt, it reasons that using a set simplifies the uniqueness check and synthesizes the correct set-based function. This case illustrates how the model combines function invocation and oracle feedback to perform inductive program synthesis. The full trace is in Section A.1.

### v2

@@ 5.5 Case Study
Figure˜4 shows an interaction trace from our benchmark. The model starts by querying the ground-truth function with edge-case inputs, aiming to probe its behavior beyond the initial examples. It then synthesizes a candidate solution using pairwise comparisons, which passes the given examples but fails on a counterexample with unhashable elements. From the error message, the model correctly infers that the ground-truth function raises a TypeError, while its own does not. On the second attempt, it reasons that using a set simplifies the uniqueness check and synthesizes the correct set-based function. This case illustrates how the model combines function invocation and oracle feedback to perform inductive program synthesis. The full trace is in Section˜A.1.
@@
6 Discussion
6.1 Results on BigCodeBench for Broader Domain Coverage
Our interactive protocol is domain-agnostic and readily extensible to any Python program. To further increase the domain coverage, we extend the original CodeARC dataset with 200 Python functions randomly sampled from BigCodeBench (Zhuo et al., 2024), which includes problems involving scientific computing (e.g., NumPy, SciPy, pandas), machine learning libraries (e.g., sklearn), visualization (e.g., matplotlib), and system libraries. This addition enhances domain diversity beyond traditional algorithm-focused problems.
Model CodeARC (Original) BigCodeBench
LLaMA-3.1-8B Base 19.3 32.5
Fine-tuned LLaMA 25.3 40.0
gpt-4o 37.8 46.5
o1-mini 53.2 40.0
o3-mini 59.5 43.0
Table 6: Success rates (%) on the original CodeARC benchmark and BigCodeBench, which covers diverse domains such as scientific computing and machine learning, extending beyond algorithm-focused problems.
Table˜6 shows that reasoning models (o3-mini, o1-mini) achieve lower success rates on BigCodeBench compared to their performance on the original CodeARC benchmarks. In contrast, non-reasoning models perform better on BigCodeBench. We attribute this to the nature of BigCodeBench tasks, which often involve domain-specific APIs and library usage, where non-reasoning models perform better. Notably, our fine-tuned model shows a substantial gain over the base model on BigCodeBench, indicating that the model fine-tuned on synthesis reasoning traces can generalize to unseen datasets.
6.2 Impact of Providing Access to a Python Interpreter
Our default evaluation restricts access to external tools such as a Python interpreter, though the CodeARC protocol can easily incorporate them via additional actions. To assess the impact, we added a code execution action, updated the prompts, and evaluated 100 sampled problems from the dataset.
Model w/o Interpreter w/ Interpreter
gpt-4o 48.5 46.0
o3-mini 69.0 71.0
Table 7: Success rates (%) with and without access to a Python interpreter.
Results shown in Table˜7 suggest that providing access to a Python interpreter does not uniformly improve model performance. For instance, while o3-mini benefits modestly (+2.0%), gpt-4o’s performance slightly decreases (-1.5%). This outcome highlights that expanding the agent’s action space may sometimes lead to suboptimal behavior.
We note that internal code simulation is a natural part of the inductive program synthesis process. Our benchmark is designed to evaluate inductive reasoning capability, and we believe that requiring models to reason without relying on external execution tools remains a meaningful and challenging setting.
@@ 7 Conclusion
Acknowledgments
We thank Yuhui Zhang and Xiaohan Wang for the discussion. This work was partially supported by a Google Research Award.
@@ A.2 Supervised Fine-Tuning Experiment Setup
A.3 Analysis of Differential Testing Tools
Program equivalence is undecidable, and no oracle can guarantee perfect correctness checking. To mitigate this, our evaluation combines two state-of-the-art differential testing tools, Pynguin (Lukasczyk & Fraser, 2022) and Mokav (Etemadi et al., 2024). This dual-oracle setup reduces the likelihood of overlooking behavioral discrepancies between the synthesized and ground-truth functions.
Empirically, the two tools produce identical outcomes on 75.4% of problems. Mokav detects additional bugs in 18.8% of cases missed by Pynguin, while the reverse holds in only 5.9%. Overall, using both tools yields a more reliable approximation of correctness in the absence of a perfect oracle.